CRY+VIS Interface Meeting
Date: 2022/3/18 13:00-13:40
Participants: Takahashi, Kimura, Miyoki, Yokozawa, Uchiyama, Hirata, Ikeda, Sato, Yasui, Ushiba, Aso, Yasui
- IX: Adjustment after the earth quake.
- EX: Adjustment after the earth quake.
- IY: Health check.
- EY: IP tuning, Health check. Go to recovery of F2.
- Replacement of flip cables: TMS.
- TM height setting: ITMY.
- Oplev: Set only for the tilt in IY.
- ACC: Installation was finished.
- Commissioning in air: Alignment of ITMY.
- Commissioning in vacuum:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Ion pumps were delivered. Setup is scheduled for the next week.
- Heater control: Test in IXV/EXV.
- TMS: Chain block should be loose. The gauge will be replaced.
Date: 2022/3/18 14:00-14:30
- Tuned the IP in EY.
- Investigated the GAS filters with the 360 camera in EY. The cable on the F2 side was very close to the pole of the EQ stop.
- Replaced the SG for the LVDT drivers in IY.
- Together with Takahashi-san, we measured the whole set of relevant transfer functions in SR2.
- We compared with measurements from July 2019 and the differences were not relevant. The behaviour of the suspension seems very stable.
- We recorded in the klog the DC amplitudes of all transfer functions for the purposes of health check.
- Today Friday I might have disturbed the the oplev.
- Later, It was possible to engage ALIGNED state with Guardian properly, so it seems fine. Still, we should measure the mirror stage transfer functions on Monday.
- The readout of the QPDs was low (a few 100s of counts per segment), so I changed he gain in the QPD boards to high with the on-board switch.
- In the BS payload, the moved the security structure screws closer to their corresponding bodies. The gaps are around 3 or 4 mm.
- The BS RM coil cables had some plastic sheets that we should have removed after installation but we forgot. We removed them today.
- Hirata-san is still considering which one of the two methods of holding the traverser limit switches is more suitable. In the method which seems more convenient, the switches assembly would be pushed with 28 kgf (one motor considered) in case the switch malfunctions.
- We recovered PR3 pitch alignment after the earthquake. We didn't have to move OSEM V1.
- In PRM oplev, I measured the distances between the optical components.
- Type-A team: Recovery of F2 in EY.
- Type-B team: PRM coil driver, Automeasurement script, Oplev in PRM
- Water close PR2 chamber.
- FC for SR2.
(Date in Kamioka)
The week from 3/21
- Takahashi: 3/22-24
- Sato: 3/22-24
- Hirata: 3/22-25
The week from 3/28
- Takahashi: 3/29-4/1
- Sato: 3/29-4/1
- Hirata: 3/29-4/1